A charge of 240C is moved when energy of 45J is applied between two points. Find the voltage between the two points.


1
Expert's answer
2021-08-16T08:39:04-0400

"E= 45 \;J \\
Q= 240 \;C \\
E= \frac{1}{2}CV^2 \\
Q=CV \\
E= \frac{1}{2}(CV)V \\
E= \frac{1}{2}QV \\
V= \frac{2E}{Q} \\
V= \frac{2 \times 45}{240}= 0.375 \;Volt"